# Bootstrapped Control Limits for Score-Based Concept Drift Control Charts

### v1

@@ 4.2 Nonlinear Oscillator System
Figure 4: Typical monitoring results using our bootstrap CL approach for the coupled nonlinear oscillator system under (a) low-noise and (b) high-noise conditions. The shifts were introduced at $i=201$ and are first detected at observations $i=225$ and $i=307$, respectively.
@@
We consider two scenarios with different noise levels, the results for which are shown in Figure 4 for a typical replicate. In the low-noise scenario (Figure 4(a); $\sigma=0.03$), the fitted model has a CV $R^{2}$ of 0.928, and the $T^{2}$ statistic increases above the CL shortly after the distribution shifts at new observation 201. The $T^{2}$ statistic remains consistently above the CL after the shift, reflecting our method’s ability to reliably detect subtle changes in the system dynamics. The high-noise scenario ($\sigma=0.3$) represents the situation that the noise is so high that the model’s predictive power becomes very poor (CV $R^{2}=0.108$). However, as seen in Fig. 4(b), our approach still successfully detects the concept drift, although there are post-shift periods over which the $T^{2}$ statistic temporarily drops below the CL.
Although the neural network is only an approximation of the underlying dynamics of Eq. (4.4) (i.e., there is some level of model misspecification), the approach is still able to detect shifts in the predictive relationship. These results align with the arguments surrounding Eq. (2.4) and in [20] that the score vector mean can capture changes in the underlying predictive relationship $\mathbb{P}(Y|\mathbf{X})$ even with model misspecification. This feature is important in practice, since fitted machine learning models are always approximations of predictive relationships.
5 Concluding Remarks

### v2

@@ 4.2 Nonlinear Oscillator System
Figure 4.3: Typical monitoring results using our bootstrap CL approach for the coupled nonlinear oscillator system under (a) low-noise and (b) high-noise conditions. The shifts were introduced at $i=201$ and are first detected at observations $i=215$ and $i=221$, respectively.
@@
We consider two scenarios with different noise levels, the results for which are shown in Figure 4.3 for a typical replicate. In the low-noise scenario (Figure 4.3(a); $\sigma=0.03$), the fitted model has a CV $R^{2}$ of 0.975, and the $T^{2}$ statistic increases above the CL shortly after the distribution shifts at new observation 201. The $T^{2}$ statistic remains consistently above the CL after the shift, reflecting our method’s ability to reliably detect subtle changes in the system dynamics. The high-noise scenario ($\sigma=0.25$) represents the situation that the noise is so high that the model’s predictive power becomes very poor (CV $R^{2}=0.137$). However, as seen in Fig. 4.3(b), our approach still successfully detects the concept drift, although there are post-shift periods over which the $T^{2}$ statistic temporarily drops below the CL.
(a) Low noise ($\sigma=0.03$, 5-fold CV $R^{2}\!\approx\!0.975$)
(b) High noise ($\sigma=0.25$, 5-fold CV $R^{2}\!\approx\!0.137$)
Figure 4.4: Empirical pointwise false-alarm rate (PFAR) for the nonlinear oscillator example under low- and high-noise conditions. The dashed horizontal line indicates the nominal level $\alpha=0.001$. Even under severe model misspecification (high noise), the bootstrap control limits maintain accurate pointwise false-alarm control prior to the change point.
Taken together, these results illustrate that the score-based monitoring approach remains effective even when the predictive model is a poor approximation to the true data-generating mechanism. Despite the neural network being only an approximation of the nonlinear dynamics in Eq. (4.4), changes in the underlying predictive relationship still induce a nonzero mean in the score vectors, leading to detectable shifts in the MEWMA statistic, as predicted by Eq. (2.4) and discussed in Zhang et al. (2023). Beyond detection behavior, Figure 4.4 further demonstrates that the proposed bootstrap control limits maintain accurate pointwise false-alarm control prior to the change point under both low- and high-noise conditions. In particular, even in the high-noise regime (CV $R^{2}\approx 0.14$), the empirical PFAR remains close to $\alpha=0.001$. This robustness is important in practice, since fitted machine learning models are invariably approximations of complex predictive relationships.
5 Conclusion